# Differential Confounding Privacy and Inverse Composition

### v3

@@ Abstract
Differential privacy (DP) has become the gold standard for privacy-preserving data analysis, but its applicability can be limited in scenarios involving complex dependencies between sensitive information and datasets. To address this, we introduce Differential Confounding Privacy (DCP), a framework that generalizes DP by accounting for broader causal relationships between secrets and datasets. DCP adopts the $(\epsilon,\delta)$-privacy framework to quantify privacy loss, particularly under the composition of multiple mechanisms accessing the same dataset. We show that while DCP mechanisms retain privacy guarantees under composition, they lack the graceful compositional properties of DP. To overcome this, we propose an Inverse Composition (IC) framework, where a leader-follower model optimally designs a privacy strategy to achieve target guarantees without relying on worst-case privacy proofs. Experimental results validate IC’s effectiveness in managing privacy budgets and ensuring rigorous privacy guarantees under composition.
@@ I Introduction
Pufferfish privacy (PP) [2] generalizes DP by extending adjacency from datasets to distributions of input datasets [3] given attacker’s prior knowledge, enabling it to incorporate domain-specific knowledge and handle data correlations more effectively. Appendix -B shows the formal definition of PP.
Both DP and PP consider the chain-rule relationship between the secret, dataset (and its distribution), and mechanism output: $S\rightarrow X\rightarrow Y$, where DP is a special case as $X$ directly contains $S$. While DP and PP provide robust privacy guarantees, they assume simplified causal structures where secrets directly identify the datasets or their distributions. However, in real-world systems such as multi-agent networks, machine learning pipelines, or distributed systems, sensitive information often arises from complex interactions and dependencies that extend beyond individual records or datasets. These systems involve derived knowledge, intermediate information, or correlations generated within internal mechanisms. In such cases, we may have to construct the relationship between $S$ and $X$ in a more general form in privacy frameworks: $S\leftrightharpoons X\rightarrow Y$. Here are a few illustrative examples (see Appendix -D for more detailed examples).
@@
To capture such general causal relationships, we extend the PP’s probabilistic framework and introduce differential confounding privacy (DCP) to quantify the privacy leakage of the secret $S$ from the output $Y$. DCP adopts the $(\epsilon,\delta)$ scheme of DP and PP to parameterize probabilistic (near) indistinguishability.
@@ I-A Related Work
The notion of DCP builds on differential privacy (DP) [1] and Pufferfish privacy (PP) [2] by generalizing the chain-rule structure between sensitive information and datasets to broader probabilistic causation relationships. A key challenge in PP frameworks is the scarcity of suitable mechanisms. Notable instantiations include Blowfish privacy [10], Wasserstein mechanisms for correlated data [11], attribute privacy [12], and mutual-information PP [13].
Composition of PP PP does not always compose [2, 11]. [2] introduce a sufficient condition, universally composable evolution scenarios (UCES), for linear self-composition, analogous to DP’s basic composition property [1]. UCES assumes a deterministic one-to-one mapping between a secret and a dataset, leveraging DP’s composition benefits by eliminating randomness in this relationship. However, in CP, where causation may flow from datasets to secrets, UCES is often infeasible. This complexity arises when another mechanism controls the state-dataset relationship, preventing the elimination of randomness.
@@ II Differential Confounding Privacy
Let $\mathcal{Q}$ be the set of such adjacent secrets. In addition, let $P_{\theta}(s)$ and $P_{\theta}(x|s)$ be the marginal and conditional densities derived from $P_{\theta}(s,x)$. In addition, let $\mathtt{D}(\cdot|\theta):\mathcal{S}\times\mathcal{S}\to\mathbb{R}$ be a metric satisfying $\mathtt{D}(s,s^{\prime}|\theta)\leq\mathtt{d}$ for a given $\mathtt{d}\geq 0$ for any $s,s^{\prime}\in\mathcal{S}$ with $P_{\theta}(s)>0$ and $P_{\theta}(s^{\prime})>0$, which defines when two secrets $s$ and $s^{\prime}$ are considered adjacent. Let $\mathcal{Q}$ be the set of such adjacent secrets.
@@
In this example, the secret $s$ is inferred information derived from $x$. Consequently, DP becomes inapplicable, as it is designed to protect individual input data records rather than properties inferred from them. Although mathematically the PP framework and the DCP appear similar, since the conditional density $P_{\theta}(x|s)$ in (2) is derived from the intrinsic probabilistic relationship between $x$ and $s$, rather than the adversary’s prior knowledge of the dataset, the standard PP framework does not directly apply to this scenario.
@@ III Characterizing the Composition
Effective Mechanism For each $\gamma_{i}:\mathcal{X}\to\Delta(\mathcal{Y}_{i})$, let $\psi_{i}:\mathcal{S}\to\Delta(\mathcal{Y}_{i})$ denote the underlying density function of $\textbf{Pr}^{s}_{\gamma}[\cdot]$ given by Equation 2. Specifically,
@@
If we have a causal chain from $S$ to $X$, DCP becomes DP. The PLRV of the composition $\mathcal{M}(\vec{\gamma})$ coincides with the PLRV of the composition of efficient mechanisms $\mathcal{N}(\vec{\psi})$. That is,
@@
Copula For any $s\in\mathcal{S}$, let $B(\cdot|s)$ denote the cumulative distribution function (CDF) and $b(\cdot|s)$ the density function of $\textbf{Pr}^{s}_{\vec{\gamma}}$ as defined in Equation 10. To simplify, we assume the outputs $y_{i}\in\mathcal{Y}_{i}$ are univariate for all $i\in[k]$. By Sklar’s theorem [29], there exists a copula $C$ for every $s\in\mathcal{S}$ such that:
@@ IV Inverse Composition
where $\textbf{L}_{\hat{b}_{s_{0}},\hat{b}_{s_{1}}}(\vec{Y})$, given by (12), is fixed but unknown. Thus, changes in $\textbf{L}^{\alpha}{s_{0},s_{1}}$ are driven by $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$, with $\alpha$ also influencing the response strategy $\pi$. Traditional forward approaches rely on generally challenging worst-case proofs, including sensitivity calculations (NP-hard [8]) required for the proper distributions of $\alpha$, and combinatorial analysis to compute $\epsilon_{g}$ and $\delta_{g}$ (#P-complete even for independent composition [6]). In contrast, IC designs $\alpha$ to elicit a follower response $\pi$ with certain properties. The key insight is that if $\alpha$ induces a $\pi$ satisfying these properties, the corresponding $\alpha$ will lead to $\textbf{L}^{\alpha}_{s_{0},s_{1}}$ that automatically achieves the desired privacy guarantees, without requiring explicit constraints on $\alpha$.
@@
For $\epsilon_{g}\geq 0$ and $\delta_{g}\in(0,1]$, define:
$\Pi[\epsilon_{g},\delta_{g}]\equiv\left\{\pi\middle|\begin{aligned} &\pi(s\mid\vec{y}_{\alpha})\geq e^{-\epsilon_{g}}P_{\theta}(s),\quad\forall s,\vec{y}_{\alpha},\\ &\mathbb{E}_{\pi}\left[\frac{\pi(S\mid\vec{y}_{\alpha})}{P_{\theta}(S)}\right]\leq\delta_{g}e^{\epsilon_{g}},\quad\forall\vec{y}_{\alpha}\end{aligned}\right\},$
@@
$\Pi[\epsilon_{g},0]\equiv\left\{\pi\middle|e^{-\epsilon_{g}}P_{\theta}(s)\leq\pi(s\mid\vec{y}_{\alpha})\leq e^{\epsilon_{g}}P_{\theta}(s),\quad\forall s,\vec{y}_{\alpha}\right\}.$
Convex Optimization Given $\epsilon_{g}\geq 0$ and $\delta_{g}\in[0,1]$, consider the following two optimization problems:
$\min_{\alpha,\pi}\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}\left[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})\right],\text{ s.t. }\pi\in\Pi[\epsilon_{g},\delta_{g}],$ (Task-1)
$\min_{\epsilon_{g},\alpha,\pi}\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}\left[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})\right],\text{ s.t. }\pi\in\Pi[\epsilon_{g},\delta_{g}].$ (Task-2)
The posterior $\mu$ is generally not convex in $\alpha$. However, for certain choices of SPSR $\mathcal{L}$, the explicit formulation of $\mu$ can be eliminated, making $\mathcal{L}$ convex in $\alpha$ and $\pi$. For example, consider ^$\hat{\mathcal{L}}(\pi,s,\vec{y}_{\alpha})=-\log(\pi(s\mid\vec{y}_{\alpha}))P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha}),$ where $P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})=\sum_{s}\alpha(y_{\alpha}\mid s)P_{\theta}(s)$ is the total probability, also the denominator of $\mu$. In this case, the expectation $\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}\left[\mathcal{L}(\pi,s,\vec{y}_{\alpha})\right]$ cancels $P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})$, reducing $\mu$ to the product of $\alpha$ and the prior $p_{\theta}$. Additionally, the set $\Pi[\epsilon_{g},\delta_{g}]$ is convex in $\pi$ and $\epsilon_{g}$ and independent of $\alpha$. Thus, by choosing $\mathcal{L}=\hat{\mathcal{L}}$, both (Task-1) and (Task-2) become convex optimization.
@@
Given $\{\theta,\mathcal{G}\}$, let $\{\mathcal{M}_{i}(\gamma_{i})\}_{i=1}^{k}$ be $k$ mechanisms, each satisfying $(\epsilon_{i},\delta_{i})$-DCP. Let $\varepsilon(\epsilon_{g},\theta)\equiv\log\left(1+\frac{e^{\epsilon_{g}}-1}{P^{*}_{\theta}}\right)$ with $P^{*}_{\theta}=\min_{s}P_{\theta}(s)>0$. When $\delta_{g}\neq 0$, let $\alpha^{*}$ be a solution to (Task-1) and $\epsilon^{*}_{g}$ a solution to (Task-2). Then, for any $\delta_{g}\in[0,1]$, the following hold:
@@
    (Task-1) For any $\epsilon_{g}\geq 0$, if $\Pi[\epsilon_{g},\delta_{g}]\neq\emptyset$, the composition $\mathcal{M}(\vec{\gamma},\gamma_{\alpha^{*}})$ is $(\varepsilon(\epsilon_{g},\theta),\delta_{g})$-DCP.
@@
    (Task-2) The composition $\mathcal{M}(\vec{\gamma})$ is $(\varepsilon(\epsilon^{*}_{g},\theta),\delta_{g})$-DCP.
@@
    When $\delta_{g}=0$, $\mathcal{M}(\vec{\gamma},\gamma_{\alpha^{\dagger}})$ is $\varepsilon(\epsilon_{g},\theta)$-DCP if $\Pi[\epsilon_{g},0]\neq\emptyset$ and $\alpha^{\dagger}$ is a solution to (Task-1). Furthermore, $\mathcal{M}(\vec{\gamma})$ is $\varepsilon(\epsilon^{\dagger}_{g},\theta)$-DCP if and only if $\epsilon^{\dagger}_{g}$ is a solution to (Task-2).
Theorem 2 establishes that the privacy strategy $\alpha$ can be designed via IC to achieve Task-1 and Task-2 by eliciting an optimal response strategy $\pi$. When $\delta_{g}\neq 0$, the optimal $\pi$ satisfying the conditions of $\Pi[\epsilon_{g},\delta_{g}]$ is sufficient for the corresponding $\alpha$ to achieve $(\epsilon_{g},\delta_{g})$-DCP. In Task-1, $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$ is constructed to ensure that $\textbf{L}^{\alpha}_{s_{0},s_{1}}$ satisfies $(\epsilon_{g},\delta_{g})$-DCP without wasting privacy budget. In Task-2, $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}=0$, and the optimal $\epsilon^{*}_{g}$ represents the tightest privacy parameter for $\textbf{L}^{\alpha}_{s_{0},s_{1}}$. When $\delta_{g}=0$, the optimal $\pi$ satisfying the conditions of $\Pi[\epsilon_{g},0]$ is sufficient for $\alpha$ to achieve $(\epsilon_{g},0)$-DCP. However, these conclusions assume an ideal scenario where the convex optimization problems can be solved exactly. Since both (Task-1) and (Task-2) are infinite-dimensional, they cannot be solved directly. The exploration of algorithmic and computational methods to address these challenges is left as future work.

### v4

@@ Abstract
Differential privacy (DP) has become the gold standard for privacy-preserving data analysis, but its applicability can be limited in scenarios involving complex dependencies between sensitive information and datasets. To address this, we introduce differential confounding privacy (DCP), a specialized form of the Pufferfish privacy (PP) framework that generalizes DP by accounting for broader relationships between sensitive information and datasets. DCP adopts the $(\epsilon,\delta)$-indistinguishability framework to quantify privacy loss. We show that while DCP mechanisms retain privacy guarantees under composition, they lack the graceful compositional properties of DP. To overcome this, we propose an Inverse Composition (IC) framework, where a leader-follower model optimally designs a privacy strategy to achieve target guarantees without relying on worst-case privacy proofs, such as sensitivity calculation. Experimental results validate IC’s effectiveness in managing privacy budgets and ensuring rigorous privacy guarantees under composition.
@@ I Introduction
Pufferfish privacy (PP) [2] generalizes differential privacy (DP) by extending the adjacency concept from datasets to distributions over datasets, conditioned on the adversary’s prior knowledge [3]. This allows PP to incorporate domain-specific assumptions and protect customizable sensitive properties of the data—extending beyond individual data points or records. Appendix -B shows the formal definition of PP.
In DP, the secret and the dataset enter the $(\epsilon,\delta)$-indistinguishability inequality through a deterministic one-to-one relationship: once the dataset is fixed, the secret (i.e., data point) is uniquely determined, and adjacent datasets correspond to distinct secret values. Hence, the randomness in the relationship between the secret and the mechanism output arises solely from the randomness of the mechanism itself. However, in real-world systems such as multi-agent networks, machine learning pipelines, or distributed systems, sensitive information often arises from complex interactions and dependencies that go beyond individual records, data attributes, or features that are observable or directly inferable from the dataset (e.g., an individual’s activity at a specific time). These systems involve derived knowledge, intermediate information, or correlations generated within internal mechanisms. Here are a few illustrative examples (see Appendix -D for more detailed examples).
@@
To capture such general relationships between $S$ and $X$, we introduce differential confounding privacy (DCP), a specialized form of the PP framework, to quantify the privacy leakage of the secret $S$ from the output $Y$. DCP adopts the $(\epsilon,\delta)$ scheme of DP and PP to parameterize probabilistic (near) indistinguishability.
@@ I-A Related Work
Differential confounding privacy (DCP) builds on the general Pufferfish privacy (PP) framework [2], which extends the deterministic, one-to-one relationship between secrets and datasets assumed in Differential privacy (DP) [1] to support more general and expressive privacy specifications. Despite its expressive power, a key challenge in the PP framework is the scarcity of practical and broadly applicable mechanisms. Notable instantiations include Blowfish privacy [10], Wasserstein mechanisms for correlated data [11], attribute privacy [12], and mutual-information-based PP [13].
Composition of PP PP does not always compose [2, 11]. [2] introduce a sufficient condition, universally composable evolution scenarios (UCES), for linear self-composition, analogous to DP’s basic composition property [1]. UCES assumes a deterministic one-to-one mapping between a secret and a dataset, leveraging DP’s composition benefits by eliminating randomness in this relationship, which corresponds to an extremely confident attacker. However, in CP, where causation may flow from datasets to secrets, UCES is often infeasible. This complexity arises when another internal mechanism controls the state-dataset relationship, preventing the elimination of randomness.
@@ II Differential Confounding Privacy
In addition, let $P_{\theta}(s)$ and $P_{\theta}(x|s)$ be the marginal and conditional densities derived from $P_{\theta}(s,x)$. In addition, let $\mathtt{D}(\cdot|\theta):\mathcal{S}\times\mathcal{S}\to\mathbb{R}$ be a metric satisfying $\mathtt{D}(s,s^{\prime}|\theta)\leq\mathtt{d}$ for a given $\mathtt{d}\geq 0$ for any $s,s^{\prime}\in\mathcal{S}$ with $P_{\theta}(s)>0$ and $P_{\theta}(s^{\prime})>0$, which defines when two secrets $s$ and $s^{\prime}$ are considered adjacent. Let $\mathcal{Q}$ be the set of such adjacent secrets.
@@
In this example, the secret $s$ is inferred information derived from $x$. Consequently, DP becomes inapplicable, as it is designed to protect individual input data records rather than properties inferred from them.
@@ III Characterizing the Composition
Effective Mechanism For each $\gamma_{i}:\mathcal{X}\to\Delta(\mathcal{Y}_{i})$, let $\psi_{i}:\mathcal{S}\to\Delta(\mathcal{Y}_{i})$ denote the underlying density function of $\textbf{Pr}^{s}_{\gamma}[\cdot]$ given by Equation (2). Specifically,
@@
If we consider a deterministic, one-to-one relationship between $S$ and $X$, DCP becomes DP. The PLRV of the composition $\mathcal{M}(\vec{\gamma})$ coincides with the PLRV of the composition of efficient mechanisms $\mathcal{N}(\vec{\psi})$. That is,
@@
Copula For any $s\in\mathcal{S}$, let $B(\cdot|s)$ denote the cumulative distribution function (CDF) and $b(\cdot|s)$ the density function of $\textbf{Pr}^{s}_{\vec{\gamma}}$ as defined in Equation (10) of Appendix -C. To simplify, we assume the outputs $y_{i}\in\mathcal{Y}_{i}$ are univariate for all $i\in[k]$. By Sklar’s theorem [29], there exists a copula $C$ for every $s\in\mathcal{S}$ such that:
@@ IV Inverse Composition
where $\textbf{L}_{\hat{b}_{s_{0}},\hat{b}_{s_{1}}}(\vec{Y})$, given by (12) of Appendix -C, is fixed but unknown. Thus, changes in $\textbf{L}^{\alpha}{s_{0},s_{1}}$ are driven by $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$, with $\alpha$ also influencing the response strategy $\pi$. Traditional forward approaches rely on generally challenging worst-case proofs, including sensitivity calculations (NP-hard [8]) required for the proper distributions of $\alpha$, and combinatorial analysis to compute $\epsilon_{g}$ and $\delta_{g}$ (#P-complete even for independent composition [6]). In contrast, IC designs $\alpha$ to elicit a follower response $\pi$ with certain properties. The key insight is that if $\alpha$ induces a $\pi$ satisfying these properties, the corresponding $\alpha$ will lead to $\textbf{L}^{\alpha}_{s_{0},s_{1}}$ that automatically achieves the desired privacy guarantees, without requiring explicit constraints on $\alpha$.
@@
For $\tau_{g}\geq 1$ and $\delta_{g}\in(0,1]$, define:
$\Pi[\tau_{g},\delta_{g}]\equiv\left\{\pi\middle|\begin{aligned} &\pi(s\mid\vec{y}_{\alpha})\geq\tau^{-1}_{g}P_{\theta}(s),\quad\forall s,\vec{y}_{\alpha},\\ &\mathbb{E}_{\pi}\left[\frac{\pi(S\mid\vec{y}_{\alpha})}{P_{\theta}(S)}\right]\leq\delta_{g}\tau_{g},\quad\forall\vec{y}_{\alpha}\end{aligned}\right\},$
@@
$\Pi[\tau_{g},0]\equiv\left\{\pi\middle|\tau^{-1}_{g}P_{\theta}(s)\leq\pi(s\mid\vec{y}_{\alpha})\leq\tau_{g}P_{\theta}(s),\quad\forall s,\vec{y}_{\alpha}\right\}.$
Convex Optimization Given $\tau_{g}\geq 1$ and $\delta_{g}\in[0,1]$, consider the following two optimization problems:
$\min_{\alpha,\pi}\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}\left[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})\right],\text{ s.t. }\pi\in\Pi[\tau_{g},\delta_{g}],$ (Task-1)
$\min_{\tau_{g},\alpha,\pi}\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}\left[\mathcal{L}(\pi,S,\vec{Y}_{\alpha})\right],\text{ s.t. }\pi\in\Pi[\tau_{g},\delta_{g}].$ (Task-2)
The posterior $\mu$ need not be convex in $\alpha$. However, if we choose an SPSR ^$\hat{\mathcal{L}}(\pi,s,\vec{y}_{\alpha})=-\log\bigl{(}\pi(s\mid\vec{y}_{\alpha})\bigr{)}\,P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})$, where $P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})=\sum_{s}\alpha(y_{\alpha}\mid s)\,P_{\theta}(s)$, then the factor $P_{\alpha,\vec{\gamma}}(\vec{y}_{\alpha})$ cancels inside the expectation $\mathbb{E}^{\mu}_{\alpha,\vec{\gamma}}[\mathcal{L}]$, leaving only the product $\alpha(y_{\alpha}\mid s)\,P_{\theta}(s)$. Moreover, $\Pi[\tau_{g},\delta_{g}]$ is convex in $\pi$ and $\tau_{g}$ and does not depend on $\alpha$. Hence, with $\mathcal{L}=\hat{\mathcal{L}}$, both (Task-1) and (Task-2) reduce to convex programs.
@@
Given $\{\theta,\mathcal{G}\}$, let $\{\mathcal{M}_{i}(\gamma_{i})\}_{i=1}^{k}$ be $k$ mechanisms, each satisfying $(\epsilon_{i},\delta_{i})$-DCP. Let $\varepsilon(\tau_{g},\theta)\equiv\log\left(1+\frac{\tau_{g}-1}{P^{*}_{\theta}}\right)$ with $P^{*}_{\theta}=\min_{s}P_{\theta}(s)>0$. When $\delta_{g}\neq 0$, let $\alpha^{*}$ be a solution to (Task-1) and $\tau^{*}_{g}$ a solution to (Task-2). Then, for any $\delta_{g}\in[0,1]$, the following hold:
@@
    (Task-1) For any $\tau_{g}\geq 1$, if $\Pi[\tau_{g},\delta_{g}]\neq\emptyset$, the composition $\mathcal{M}(\vec{\gamma},\gamma_{\alpha^{*}})$ is $(\varepsilon(\tau_{g},\theta),\delta_{g})$-DCP.
@@
    (Task-2) The composition $\mathcal{M}(\vec{\gamma})$ is $(\varepsilon(\tau^{*}_{g},\theta),\delta_{g})$-DCP.
@@
    When $\delta_{g}=0$, $\mathcal{M}(\vec{\gamma},\gamma_{\alpha^{\dagger}})$ is $\varepsilon(\tau_{g},\theta)$-DCP if $\Pi[\tau_{g},0]\neq\emptyset$, for $\tau_{g}\geq 1$, and $\alpha^{\dagger}$ is a solution to (Task-1). Furthermore, $\mathcal{M}(\vec{\gamma})$ is $\varepsilon(\tau^{\dagger}_{g},\theta)$-DCP if and only if $\tau^{\dagger}_{g}$ is a solution to (Task-2).
Theorem 2 establishes that the privacy strategy $\alpha$ can be designed via IC to achieve Task-1 and Task-2 by eliciting an optimal response strategy $\pi$. When $\delta_{g}\neq 0$, the optimal $\pi$ satisfying the conditions of $\Pi[\tau_{g},\delta_{g}]$ is sufficient for the corresponding $\alpha$ to achieve $(\varepsilon(\tau_{g},\theta),\delta_{g})$-DCP. In Task-1, $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}$ is constructed to ensure that $\textbf{L}^{\alpha}_{s_{0},s_{1}}$ satisfies $(\varepsilon(\tau_{g},\theta),\delta_{g})$-DCP without wasting privacy budget. In Task-2, $\textbf{L}_{\alpha_{s_{0}},\alpha_{s_{1}}}=0$, and the optimal $\epsilon^{*}_{g}=\varepsilon(\tau^{*}_{g},\theta)$ represents the tightest privacy parameter for $\textbf{L}^{\alpha}_{s_{0},s_{1}}$. When $\delta_{g}=0$, the optimal $\pi$ satisfying the conditions of $\Pi[\tau_{g},0]$ is sufficient for $\alpha$ to achieve $(\varepsilon(\tau_{g},\theta),0)$-DCP. However, these conclusions assume an ideal scenario where the convex optimization problems can be solved exactly. Since both (Task-1) and (Task-2) are infinite-dimensional, they cannot be solved directly. The exploration of algorithmic and computational methods to address these challenges is left as future work.